# An Appropriate Probability Model for the Bell Experiment

### v2

@@ Abstract
The Bell inequality constrains the outcomes of measurements on pairs of distant entangled particles. The Bell contradiction states that the Bell inequality is inconsistent with the calculated outcomes of these quantum experiments. This contradiction led many to question the underlying assumptions, viz. so-called realism and locality. The probability model underlying the Bell inequality is generally left implicit. This implicit consensus model assumes four simultaneously observable detector settings. The Bell contradiction follows from this assumption.
We propose an explicit probability model for the CHSH version of the Bell experiment. This model has only two simultaneously observable detector settings per measurement, and therefore does not assume realism. The quantum expectation now becomes a conditional expectation, given the two detector settings. This probability model is in full agreement with both quantum mechanics and experiments. In this model the notion of Bell contradiction has no meaning. Furthermore, the proposed probability model is statistically local, and is not Bell-separable. The latter implies that either hidden variables must be ruled out, or that locality must be violated, in agreement with Bell’s conclusion.
@@ 1 Introduction
J.S. Bell describes an experiment [1] in which an electron and a positron, formed somehow in a singlet spin state, move freely in opposite directions. This experiment was inspired by the EPR paradox [2]. He derives an inequality, labeled (15) in [1], that constrains the expected outcomes of the measurements of selected components of the spins of these two particles. Today this inequality is known as the Bell inequality.
The Bell inequality connects two different formalisms: quantum mechanics and probability models. In quantum mechanics the state of a system is modeled as a vector in a Hilbert space. The evolution of this state is governed by a unitary operator and is deterministic. An observation (measurement) of this state is by the application of a Hermitean operator which yields a probability value for each eigenvector of that operator. This measurement procedure is inherently probabilistic (non-deterministic) and thus requires an appropriate probability model.
In the body of this paper we focus on the more general CHSH version [3] of the Bell experiment with four detector settings, two detector on each site. Furthermore, we assume that the particles are entangled photons, and that the detectors are polarimeters. [^2] The main findings of our work can straightforwardly be adapted to the original Bell experiment with three detector settings and to fermions.
The spins of the two prepared particles (photons) are measured by detectors (polarimeters) $A$ and $B$. Each detector has two different settings, and the selection among the available detector settings is possibly random, and possibly after the photon pair has been prepared. Each individual measurement outcome has either value +1 (spin up) or -1 (spin down). The corresponding stochastic variables for the detector settings of detector $A$ are denoted by $X_{0}$, $X_{1}$, and for detector $B$ by $Y_{0}$, and $Y_{1}$. The CHSH version [3] of the Bell inequality is:
@@
Next, Bell also calculates the outcomes of these experiments, by applying the relevant observables (self-adjoint operators) to the quantum state of the photon pair. He thus arrives at the Bell contradiction: these calculated outcomes contradict the inequality. The Bell inequality also inspired numerous experiments, including [4] and [5]. These experiments typically demonstrate “loophole-free Bell-inequality violation”.
@@
    The setting of detector A does not impact the outcome Y on the other detector and vice versa. This assumption is known as (statistical) locality.
The details of these assumptions are still being widely debated [^3] To date, [1] has been cited over 16 thousand times. among physicists, probability theorists, and philosophers, see for example [7, 8, 9, 10].
In this paper we offer a different and fresh perspective on the Bell contradiction. Rather than questioning the underlying assumptions, we question the underlying probability model itself. A probability model comprises an outcome space, an event space and a probability measure. These notions will be reviewed in some detail in Section 2. Interestingly and significantly, an appropriate outcome space for the Bell experiment is generally left implicit. The few exceptions include [11, 12, 13], where the experiment is modeled as a Bayesian network. However, these Bayesian-network models are inherently factorizable with conditional probabilities, and therefore disagree with quantum mechanics. In Section 3, we derive the quantum mechanical description of the Bell experiment. In Section 4 we review the model with four observables that seems to have been used implicitly by other reseachers and that leads to the Bell paradox. Bell [1] derived his inequality in a similar way. Next, in Section 5, we present an appropriate probability space for the Bell experiment, including a probability measure, derived by quantum mechanics in the former section, and calculate the expected values of the observables in the Bell inequality. In this model there is no Bell contradiction.
Bell introduced a so-called hidden variable to see if it could account for the observed correlations among measurement outcomes. Technically, this means that the underlying probability measure must be separable in a well-defined sense. In Section 6 we analyze various properties of the proposed probability model, and prove that the proposed probability measure is not Bell-separable. The implication is that either local hidden-variable theories must be ruled out, or that the underlying assumption of locality is violated.
In Appendix A we give an alternative proof of non-separability and we also show that if separability would hold, then the system should be deterministic.
@@ 3 A quantum mechanical model
This expectation is in agreement with those of [3] and [4]. [^4] The minus sign preceding the $\cos()$ is sometimes omitted due to different conventions. Figure 3 in [4] shows an excellent agreement between the measured data and (3.9).
@@ 4 A model with four observables
Note that it $\big|\mathbb{E}[X_{\alpha}Y_{\beta}\big|\leq 1$ and so, for any probability measure on the outcome space of Table 1 (4.4) holds. Using (3.9) it follows that
@@
Next, we will use the specific values for the detectors $\alpha:a_{0},a_{1}$ and two for $\beta:b_{0},b_{1}$. (Recall that these settings correspond to the random variables $X_{0},X_{1},Y_{0},Y_{1}$ as used in Table 1.) Following [6],
@@ 5 A model with 2 observables and 2 detector settings
An appropriate probability model for the Bell experiment can be based on a different set of four stochastic variables: the two observables $X$ and $Y$ (the two measurements) and the two settings $A$ and $B$. Furthermore, the quantum mechanical expectation $\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}$ now becomes the conditional expectation
@@
that is, the expectation to observe $X$ and $Y$ given detector settings $A$ and $B$. The corresponding (conditional) probability space is shown in Table 2. Note that the outcomes of $A$ and $B$ must be known in order to compute the simultaneous probability of $(X,Y)$. The tabulated probabilities are based on (3.8), and hence agree with the quantum mechanical model and its expermental validation [4].
@@
So, this is consistent with (4.7) and (4.8) and therefore we don’t have a contradiction. The step from (4.3) to (4.4) is allowed for expectations, but is incorrect for conditional expectations under different conditions. This is the cause of the earlier contradiction. [^5] Summing conditional expectations looks like a simple and innocent operation: $\mathbb{E}[X+Y\!\;|\;\!H_{1}]=\mathbb{E}[X\!\;|\;\!H_{2}]+\mathbb{E}[Y\!\;|\;\!H_{3}]$ , is only true if $H_{1}=H_{2}=H_{3}$. As an example consider the sum of dice $X$ and $Y$. With $\mathbb{E}[X]=\mathbb{E}[Y]=3.5$ it follows that $\mathbb{E}[X+Y]=7$. However, $\mathbb{E}[X\!+\!Y\;|\;X\!=\!Y]~=~\mathbb{E}[X\!+\!Y\;|\;X\!\neq\!Y]~=~7$. Therefore, $\mathbb{E}[X\!+\!Y]~\neq~\mathbb{E}[X\!+\!Y\;|\;X\!=\!Y]+\mathbb{E}[X\!+\!Y\;|\;X\!\neq\!Y]$.
@@
$\begin{array}[]{lll}\mathbb{E}[\tilde{X}_{\alpha}\tilde{Y}_{\beta}]&=&\sum_{x,y}x.y.\mathbb{P}[X\!=\!x,Y\!=\!y,A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\sum_{x,y}x.y.\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta].\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\mathbb{E}[XY\;|\;A\!=\!\alpha,B\!=\!\beta].\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}.\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\end{array}$ (5.4)
@@
because it is more natural to take averages over the outcomes per setting. This suggests that the experimental outcomes of $\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}$ correspond to conditional expectations, which explains the contradiction.
6 Some consequences
With the appropriate probability model of Section 5 we got rid of the contradiction. Here we consider two properties of the model: statistical locality and (non-)separability.
@@ 6.1 Statistical locality
From Table 3 it can be observed that $\mathbb{P}[X\!=\!x\;|\;A=\alpha,B\!=\!\beta]=\frac{1}{2}$. So, the outcome of detector A does not depend on the setting of detector B, and vice versa. This property is known as (statistical) locality [9] and =also as no-signalling [8]. More precisely, probability measure $\mathbb{P}$ is statistically local if
$\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,B\!=\!\beta]~=~\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha]~,$ (6.1)
and similarly for $Y$. From Table 3 we easily derive that $\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,B\!=\!\beta]=\frac{1}{2}$. Hence $\mathbb{P}[X\!=\!x|A\!=\!\alpha]=\frac{1}{2}$ as well. Note that this property holds for all detector settings and for any probability $\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]$, so also for any arrangement made by the observers, not withstanding that the particles are entangled. So, if an observer of one particle knows the outcome, he also knows what the other observer will see.
6.2 Non-separability
Following [1], we introduce a hidden variable in the probability model. In order to accommodate this hidden variable, the outcome space for the Bell experiment is extended with a random variable $L$, where the range of $L$ is $\Lambda$ 666which is assumed here to be at most countable..
Let random variable $L$ denote the hidden variable with $L(\omega)=\lambda$ if $\omega=(x,y,\alpha,\beta,\lambda)$ and probability distribution $\mathbb{P}[L=\lambda]=\rho(\lambda)$. This model containing a hidden variable is separable [1], [9] if, for all $x,y,\alpha,\beta$ :
$p(x,y|\alpha,\beta):=\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta]=\sum_{\lambda\in\Lambda}p_{1}(x|\alpha,\lambda).p_{2}(y|\beta,\lambda).\rho(\lambda)~,$ (6.3)
where
$\begin{array}[]{lll}p_{1}(x|\alpha,\lambda)&:=&\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,L=\lambda]\\[2.84526pt] p_{2}(y|\beta,\lambda)&:=&\mathbb{P}[Y\!=\!y\;|\;B\!=\!\beta,L=\lambda]~.\end{array}$ (6.4)
Of course, the distribution should comply with the quantum-mechanical model, see Table 2. Critically, $L$ is independent of the detector settings $A$ and $B$. This assumption is known as statistical independence or measurement independence. See for example [9] for more detail.

### v3

# arXiv:2302.05174v3[quant-ph] 01 Mar 2026

@@ Abstract
The Bell inequality constrains the outcomes of measurements on pairs of distant entangled particles. The Bell contradiction states that the Bell inequality is inconsistent with the calculated outcomes of these quantum experiments. This contradiction led many to question the underlying assumptions, viz. so-called realism and locality. The probability model underlying the Bell inequality is generally left implicit. We propose an explicit probability model for the CHSH version of the Bell experiment. This model has only two simultaneously observable detector settings per measurement, and therefore does not assume realism. The quantum expectation now becomes a conditional expectation, given the two detector settings. This probability model is in full agreement with both quantum mechanics and experiments. As a result, the model satisfies the Bell inequality; there are no so-called violations. We extend this model to include a hidden variable. This extended model is not Bell-separable. This non-separability implies that the model is non-deterministic or non-local (or both).
@@ 1 Introduction
J.S. Bell describes an experiment [1] in which “a pair of spin one-half particles, formed somehow in a singlet spin state and moving freely in opposite directions”. This experiment was inspired by the EPR paradox [2]. He derives an inequality, labeled (15) in [1], that constrains the expected outcomes of the measurements of selected components of the spins of these two particles. Today this inequality is known as the Bell inequality.
The Bell inequality connects two different formalisms: quantum mechanics and probability models. In quantum mechanics the state of a system is modeled as a vector in a Hilbert space. The evolution of this state is governed by a unitary operator and is deterministic. An observation (measurement) of this state is by the application of a Hermitean operator which yields a probability value for each eigenvector of that operator. This measurement procedure is inherently probabilistic (non-deterministic) and thus requires a (formal) probability model.
In the body of this paper we focus on the more general CHSH version [3] of the Bell experiment with four detector settings, two detector on each site. The spins of the two prepared particles (photons) are measured by detectors (polarimeters) $A$ and $B$. (The main findings of our work can straightforwardly be adapted to the original Bell experiment with three detector settings and to fermions.) Each detector has two different settings, and the selection among the available detector settings is possibly random, and possibly after the photon pair has been prepared. Each individual measurement outcome has either value +1 (spin up) or -1 (spin down). The corresponding stochastic variables for the detector settings of detector $A$ are denoted by $X_{0}$, $X_{1}$, and for detector $B$ by $Y_{0}$, and $Y_{1}$. The CHSH version [3] of the Bell inequality is:
@@
Next, Bell also calculates the outcomes of these experiments, by applying the relevant observables (self-adjoint operators) to the quantum state of the photon pair. He thus arrives at the Bell contradiction: these calculated outcomes contradict the inequality. The Bell inequality also inspired numerous experiments, including [4, 5]. These experiments typically demonstrate “loophole-free Bell-inequality violation”.
@@
    The setting of detector A does not impact the outcome Y on the other detector and vice versa. This assumption is known as (statistical) locality, see Section 6.1.
The details of these assumptions are still being widely debated among physicists, probability theorists, and philosophers, see for example [7, 8, 9]. To date, Ref. [1] has been cited over 16 thousand times.
In this paper we offer a different and fresh perspective on the Bell contradiction. Rather than questioning the underlying assumptions, we question the underlying probability model itself. A probability model comprises an outcome space, an event space and a probability measure. These notions will be reviewed in some detail in Section 2. Interestingly and significantly, a suitable outcome space for the Bell experiment is generally left implicit. The few exceptions include [5, 10, 11], where the experiment is modeled as a Bayesian network. However, these Bayesian-network models are inherently factorizable with conditional probabilities, and therefore disagree with quantum mechanics. In Section 3, we derive the quantum mechanical description of the Bell experiment. In Section 4 we review the model with four observables that seems to have been used implicitly by other researchers and that leads to the Bell paradox. Bell [1] derived his inequality in a similar way. Next, in Section 5, we present a suitable probability space for the Bell experiment, including a probability measure, derived by quantum mechanics in the former section, and calculate the expected values of the observables in the Bell inequality. In this model there is no Bell contradiction.
Bell introduced a so-called hidden variable to see if it could account for the observed correlations among measurement outcomes. In Section 6 we extend the proposed model to include such a hidden variable. The probability measure of this extended model is shown non-separable in the Bell-sense. Furthermore, this non-separability implies that the model is non-local, or non-deterministic, or both. Appendix A offers an alternative proof of this non-separability.
@@ 3 A Quantum Mechanical Model
In a quantum mechanical model of the Bell experiment the quantum state of an entangled photon pair can be described by
Here ${\bf{e}}_{0}=(1,0)^{\top}$ and ${\bf{e}}_{1}=(0,1)^{\top}$ the two base spin vectors, also known as up and down.
The spin of a photon can be detected by a polarimeter, which is modeled by a special type of self-adjoint operator (Hermitean matrix) $F_{\!\theta}:\mathbb{C}^{2}\to\mathbb{C}^{2}$,
$F_{\theta}=\begin{bmatrix}\cos(2\theta)&\sin(2\theta)\\ \sin(2\theta)&-\cos(2\theta)\end{bmatrix}~$, (3.5)
where $\theta\in[0,\pi)$ is the angle of the polarizer with respect to some reference frame perpendicular to the axis of the experiment. $F_{\theta}$ is the difference of two orthogonal projections
$F_{\theta}={\bf{f}}_{\theta}{\bf{f}}_{\theta}^{\top}-\tilde{{\bf{f}}}_{\theta}\tilde{{\bf{f}}}_{\theta}^{\top}\!,~~~{\bf{f}}_{\theta}=\begin{bmatrix}\cos\theta\\ \sin\theta\end{bmatrix},~~\tilde{{\bf{f}}}_{\theta}={\bf{f}}_{\theta+\frac{\pi}{2}}=\begin{bmatrix}-\sin\theta\\ ~\cos\theta\end{bmatrix}~$. (3.6)
The outcomes (eigenvalues) of a detection are +1+1 and 1-1, with respective eigenvectors ${\bf{f}}_{\theta}$ and $\tilde{{\bf{f}}}_{\theta}$. Note that this covers all cases, because ${\bf{f}}_{\theta\,\pm\,\pi}=-{\bf{f}}_{\theta}$.
@@
This expectation is in agreement with those of [3, 4]. (The minus sign preceding the $\cos()$ is sometimes omitted due to different conventions.) Figure 3 in [4] shows an excellent agreement between the measured data and (3.12).
@@ 4 A Model with Four Observables
Note that it $\big|\mathbb{E}[X_{\alpha}Y_{\beta}\big|\leq 1$ and so, for any probability measure on the outcome space of Table 1 Equation (4.4) holds. Using (3.12) it follows that
@@
Next, we will use the specific values for the detectors $\alpha:a_{0},a_{1}$ and two for $\beta:b_{0},b_{1}$. (Recall that these settings correspond to the random variables $X_{0},X_{1},Y_{0},Y_{1}$ as used in Table 1). Following [6],
@@ 5 A Model with 2 Observables and 2 Detector Settings
A suitable probability model for the Bell experiment can be based on a different set of four stochastic variables: the two observables $X$ and $Y$ (the two measurements) and the two settings $A$ and $B$. Furthermore, the quantum mechanical expectation $\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}$ now becomes the conditional expectation
@@
that is, the expectation to observe $X$ and $Y$ given detector settings $A$ and $B$. The corresponding (conditional) probability space is shown in Table 2. Note that the outcomes of $A$ and $B$ must be known in order to compute the simultaneous probability of $(X,Y)$. The tabulated probabilities are based on (3.11), and hence agree with the quantum mechanical model and its experimental validation [4].
@@
From Table 2 it can be observed that
$\begin{array}[]{lllll}\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,B\!=\!\beta]&=&\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha]&=&\frac{1}{2}\\[2.84526pt] \mathbb{P}[Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta]&=&\mathbb{P}[Y\!=\!y\;|\;B\!=\!\beta]&=&\frac{1}{2}~.\end{array}$ (5.2)
Clearly, the outcome $X$ does not depend on the setting of detectors $A$ and $B$. Note that this property holds for all detector settings and for any probability $\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]$.
@@
So, this is consistent with (4.7) and (4.8) and therefore we don’t have a contradiction. The step from (4.3) to (4.4) is allowed for expectations, but is incorrect for conditional expectations under different conditions. This is the cause of the earlier contradiction. Aside: conditional expectations looks like a simple and innocent operation: $\mathbb{E}[X+Y\!\;|\;\!H_{1}]=\mathbb{E}[X\!\;|\;\!H_{2}]+\mathbb{E}[Y\!\;|\;\!H_{3}]$, is only true if $H_{1}=H_{2}=H_{3}$. As an example consider the sum of dice $X$ and $Y$. With $\mathbb{E}[X]=\mathbb{E}[Y]=3.5$ it follows that $\mathbb{E}[X+Y]=7$. However, $\mathbb{E}[X\!+\!Y\;|\;X\!=\!Y]~=~\mathbb{E}[X\!+\!Y\;|\;X\!\neq\!Y]~=~7$. Therefore, $\mathbb{E}[X\!+\!Y]~\neq~\mathbb{E}[X\!+\!Y\;|\;X\!=\!Y]+\mathbb{E}[X\!+\!Y\;|\;X\!\neq\!Y]$.
@@
$\begin{array}[]{lll}\mathbb{E}[\tilde{X}_{\alpha}\tilde{Y}_{\beta}]&=&\sum_{x,y}x\cdot y\cdot\mathbb{P}[X\!=\!x,Y\!=\!y,A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\sum_{x,y}x\cdot y\cdot\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta]\cdot\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\mathbb{E}[XY\;|\;A\!=\!\alpha,B\!=\!\beta]\cdot\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\\[1.42262pt] &=&\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}\cdot\mathbb{P}[A\!=\!\alpha,B\!=\!\beta]\end{array}$ (5.5)
@@
because it is more natural to take averages over the outcomes per setting. This suggests that the experimental outcomes of $\boldsymbol{\big\langle}X_{\alpha}\otimes Y_{\beta}\boldsymbol{\big\rangle}$ correspond to conditional expectations, which explains the contradiction. Below this model will be referred to as the 2-observables model.
6 The 2-Observables Model Extended with Hidden Variables
Following [1], we extend the proposed probability model with a hidden variable. In order to accommodate this hidden variable, the outcome space for the Bell experiment is extended with a random variable $\in\Lambda$, which is assumed here to be at most countable.
Let random variable $L$ denote the hidden variable with $L(\omega)\!=\!\lambda$ for $\omega\!=\!(x,y,\alpha,\beta,\lambda)$. The probability measure of this extended model must be consistent with Table 2 and hence must satisfy
$\begin{array}[]{lll}&&\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!a,B\!=\!b]\\[2.84526pt] &=&\sum_{\lambda\in\Lambda}\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!a,B\!=\!b,L\!=\!\lambda]\cdot\mathbb{P}[A\!=\!a,B\!=\!b,L\!=\!\lambda]~.\end{array}$ (6.2)
Critically, $L$ must be independent of the detector settings $A$ and $B$. This assumption is known as statistical independence or measurement independence [9]:
$\mathbb{P}[A\!=\!a,B\!=\!b,L\!=\!\lambda]~=~\mathbb{P}[A\!=\!a,B\!=\!b]\cdot\mathbb{P}[L\!=\!\lambda]~.$ (6.3)
This assumption is consistent with the idea that $\lambda$ is given before $A$ and $B$ are selected. In this section we prove:
  1. 1.
    This extended probability model is non-separable.
  2. 2.
    The non-separability of the model implies that it is non-deterministic or non-local (or both).
These properties are according to established definitions.
@@ 6.1 Statistical Locality
With (5.2) we observed that the outcome $X$ does not depend on the setting of detectors $A$ and $B$, and likewise for outcome $Y$. The property of (statistical) locality [9] is defined by
$\begin{array}[]{lll}\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,B\!=\!\beta,L\!=\!\lambda]&=&\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,L\!=\!\lambda]\\[2.84526pt] \mathbb{P}[Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta,L\!=\!\lambda]&=&\mathbb{P}[Y\!=\!y\;|\;B\!=\!\beta,L\!=\!\lambda]~.\end{array}$ (6.4)
Note that this must hold for all $\alpha,\beta,\lambda$. Paraphrasing [9]: “Thus, an observer cannot distinguish, via any local measurement based on detector setting $A$, whether a distant observer has carried out measurement using detector setting $B$ or $B^{\prime}$), even given knowledge of the underlying variable $\lambda$ . This property, also known as parameter independence and as no-signalling [8], is justified by the principle of relativity when the measurement subprocedures are carried out in spacelike separated regions [1].” Statistical locality is not assumed a priori to be a property for the extended model including hidden variable.
6.2 Bell Separability
The probability model including a hidden variable, is separable [1, 9] if, for all $x,y,\alpha,\beta$:
$\begin{array}[]{lll}p(x,y|\alpha,\beta)&=&\sum_{\lambda\in\Lambda}p_{1}(x|\alpha,\lambda)\cdot p_{2}(y|\beta,\lambda)\cdot\rho(\lambda)\\[2.84526pt] p_{1}(x|\alpha,\lambda)&:=&\mathbb{P}[X\!=\!x\;|\;A\!=\!\alpha,L\!=\!\lambda]\\[2.84526pt] p_{2}(y|\beta,\lambda)&:=&\mathbb{P}[Y\!=\!y\;|\;B\!=\!\beta,L\!=\!\lambda]~,\end{array}$ (6.5)
where $p(x,y|\alpha,\beta)$ is used as a shorthand for $\mathbb{P}[X\!=\!x,Y\!=\!y\;|\;A\!=\!\alpha,B\!=\!\beta]$, and $\rho(\lambda)$ is a shorthand for $\mathbb{P}[L=\lambda]$. The extended probability measure is repeated in Table 4, including an assumed factorization given hidden variable $\lambda$.
$a$ $a$ AA
$b$ $b$ $B$
$X$ $Y$ $\lambda$ $L$
1 1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $p(a,\lambda)\cdot(1-q(b,\lambda))$
-1 1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $(1-p(a,\lambda))\cdot(1-q(b,\lambda))$
1 -1 $\frac{1}{2}\cos^{2}(a\!-\!b)$ $p(a,\lambda)\cdot q(b,\lambda)$
-1 -1 $\frac{1}{2}\sin^{2}(a\!-\!b)$ $(1-p(a,\lambda))\cdot q(b,\lambda)$
Separability, for $\alpha=\beta$ would imply
$\begin{array}[]{lll}\sum_{\lambda\in\Lambda}~p_{1}(1|\alpha,\lambda)\cdot p_{2}(1|\alpha,\lambda)\Big)\rho(\lambda)&=&0\\ \sum_{\lambda\in\Lambda}~(1-p_{1}(1|\alpha,\lambda))\cdot(1-p_{2}(1|\alpha,\lambda))\rho({\lambda})&=&0~.\end{array}$ (6.6)
Hence, $1-p_{1}(1|\alpha,\lambda)-p_{2}(1|\alpha,\lambda)=0$ for all $\lambda\!\in\!\Lambda$, with $\rho(\lambda)\!>\!0$. Substitution of $p_{2}(1|\alpha,\lambda)=1-p_{1}(1|\alpha,\lambda)$ in (6.6) gives
$\sum_{\lambda\in\Lambda}p_{1}(1|\alpha,\lambda)\cdot\Big(1-p_{1}(1|\alpha,\lambda)\Big)\rho(\lambda)=0~.$ (6.7)
This constraint on $p_{1}$ can be rephrased as
$p_{1}(x|\alpha,\lambda)\in\{0,1\}\hskip 56.9055ptp_{2}(y|\beta,\lambda)\in\{0,1\}~.$ (6.8)
This property relates to (non-)determinism, which is the topic of Section 6.3.